OpenStudy (anonymous):

√6(4√12+5√8) i really need help!!!!! I do not know the steps to take to get the answer.

OpenStudy (anonymous):

Distribute \[\sqrt{6}\] which multiplies with the other roots

OpenStudy (anonymous):

you then get 4√72+5√48

OpenStudy (anonymous):

Simplify 4√72+5√48 =4√36*√2 + 5 √16*√3 Square root of 36 equals 6 Square root of 16 equals 4 So, 4*6√2 +5*4√3 =24√2 + 20√3 You can't add or simplify any further so that's the final answer.